If Nicolai earns an 8 percent after-tax rate of return, $16,500 today would be worth how much to Nicolai in five years? Use future value of $1. (Round discount factor(s) to four decimal places.)

$15,279.

$11,238.

$17,820.

$16,500.

None of the choices are correct.
Answer #1

Answer : None of the choices are correct

future value = Present value (1 + r)n

future value =16500 (1 + 0.08 )

future value =16500(1.08)5

future value =16500 (1.469)

future value =24239 (Approx)